OpenStudy (anonymous):

When a coin is flipped, there are two possible outcomes: heads or tails. If a coin is flipped three times, what is the probability that there will be tails all three times?

OpenStudy (anonymous):

I can post the possible answers if necessary

OpenStudy (anonymous):

Don't post possible answers.

OpenStudy (anonymous):

OK

OpenStudy (kropot72):

The 3 flips are independent trials each with a probability of 1/2 for tails. Probability of 3 tails in a row is \[\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}=you\ can\ calculate\]

OpenStudy (mathstudent55):

Figure it out yourself the same we did the previous question. Write a table with all possible outcomes. Then see what fraction your desired outcome is of the total number of outcomes.

OpenStudy (anonymous):

0.125

OpenStudy (anonymous):

Thanks!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (mathstudent55):

 There are 8 possible outcomes, and only 1 is three tails, so the probabilty of three tails is 1/8